OpenStudy (anonymous):

Write the expression in standard form. 5/3-15i

OpenStudy (anonymous):

you'll want to multiply by the conjugate of the denominator. notice we are basically multiplying by 1. \[\frac{ 5 }{ 3 - 15i } = \frac{ 5 }{ 3 - 15i }\frac{ 3 + 15i }{ 3 + 15i }\] now we got i out of the denominator (since we're left with constants and i^2) it's ok to have i in the numerator. solve/expand like the last problem :)

OpenStudy (jim766):

you good with this?

OpenStudy (anonymous):

I'm not sure yet

OpenStudy (anonymous):

My answers are: 5/78-25/78i 5/78+25/78i -5/78+25/78i -5/78-25/78i

OpenStudy (anonymous):

I am so not good with this.

OpenStudy (jim766):

those are the answers you got, or the answer choices they gave you?

OpenStudy (anonymous):

The ones they gave me.

OpenStudy (jim766):

I think the i's are supposed to go with the top numbers...

OpenStudy (jim766):

\[\frac{ 5 }{ 3-15i }\frac{ 3+ 15i }{ 3 + 15 i }\] \[\frac{ 15+ 75i }{ 9-225(-1) }\]

OpenStudy (anonymous):

They are just out to the side of the fraction in the problem

OpenStudy (jim766):

do you see where I got this?

OpenStudy (anonymous):

Yes

OpenStudy (jim766):

\[\frac{ 15 + 75i }{ 9+225 }\] \[\frac{ 15 + 75i }{ 234 }\] \[\frac{ 15 }{ 234 }+ \frac{ 75i }{ 234 }\] how about this

OpenStudy (anonymous):

Yes

OpenStudy (jim766):

\[\frac{ 15 }{ 234 }+\frac{ 75 }{ 234 }i\] \[\frac{ 5 }{ 78 }+ \frac{ 25 }{ 78 }i\] and now this?

OpenStudy (anonymous):

No, you lost me.

OpenStudy (jim766):

I just reduced the fractions... What number divides evenly in both 15 and 234? 3 15 div 3 5 --- 234 div 3 78

OpenStudy (jim766):

75 div 3 25 -- --- 234 div 3 78

OpenStudy (anonymous):

Okay, I've got you now. Thank you so much!

OpenStudy (jim766):

yw